Question 1 Not yet answered Marked out of 25,00 P Flag question Suppose that the sales function for a product A has been estimated as Q7-1.6PL 2Pct.8Ps+.0005Y+.0004A Where, QL is the quantity of product L demanded, in thousand per month PL is the price of product L, which is currently $15 Pc is the price of product C, which is currently $25 Ps is the price of product S, which is currently $20; Y is the level of per capita income, which is currently $20,000; and A is monthly advertising, which is currently $15,000. Calculate expected sales of L at current values for all variables in the sales function for L. Select one: O a. 10 thousand o b. 12 thousand O c. 1 5 thousand o d. 8 thousand Next page 4 Financial Statements Jump to... Quiz navigation Finish attermpt
Answer #1

We have the sales equation which is QL=7-1.6PL-0.2PC +0.8Ps+ 0.0005Y + 0.0004A

Substituting the values given in this equation, we get

QL= 7-(1.6*15)-(0.2*25)+(0.8*20)+(0.0005*20000) +(0.0004*15000)

QL = 7-24-5+16+10+6

QL = 39-29

QL = 10

So the expected sales of L at current values is 10 thousand.

So option a is the correct answer.